OpenStudy (anonymous):

Let f(x) = (x2 + 3x – 4) and g(x) = (x + 4).

OpenStudy (kainui):

I'll let them.

OpenStudy (anonymous):

so??

OpenStudy (hba):

You have to find f(g(x)= Right ?

OpenStudy (anonymous):

yes but how do you do that?

OpenStudy (hba):

You just plug in g(x) which is x+4 in f(x)= (x2 + 3x – 4) f(g(x))=f(x+4)= Just put x+4 in place of x got it ?

OpenStudy (anonymous):

not really...

OpenStudy (hba):

f(g(x)=f(x+4)=[(x+4)^2+3(x+4)-4] Now solve it more.

hero (hero):

Don't tell me you're still working on it. Did you figure it out yet @mathhelp9 ?

OpenStudy (anonymous):

sorry hero. it was very late i had to sleep. thanks for your help. i think i got it!

OpenStudy (hba):

Ok good :)

hero (hero):

If you figured it out, you're supposed to post what you got just in case you made a mistake so that we can confirm your correct answer.

OpenStudy (anonymous):

i was totally off track. i couldn't finish it.

OpenStudy (hba):

Show us your working.... f(g(x)=f(x+4)=[(x+4)^2+3(x+4)-4]

OpenStudy (anonymous):

Ok. This is what I got. Pls verify if this is right. thx for your helo. f(x) • g(x) = f(g(x)) f(x+4) = (x+4)^2 + 3 (x+4)-4 f(x+4)=x^2+11x+24 If x is positive: (x+3)(x+8)=f(x+4)